RFBG Commissioning with Adaptive Secondary - 05 April 2010 UT

Night 17:

Observer: DMiller FBian JHill (Tucson)
Telescope Operator: JHill (Tucson)
Telescope Support: MMidkiff
Adaptive Operator & Wind Monitor: RBriguglio (LBTO)
Vibration Operator: JEsguerra

Summary

More training of Doug (and Fuyan) to operate the telescope.

Collected more transformation data with seeing 0.8 - 1.2 arcsec. Residuals seem worse than previously seen.

Had various problems with the shell RIPing or resetting.

Collected Field Aberration Data with the reversed Tip setting from last night.

Problems with the Right Instrument Cooling loop. See Issue 2648.

Completed the DX M3 tests as the seeing degrades beyond 1.5 arcsec.

Had problems with a DX primary mirror panic at the end of the night. See Issue 2649.

Preparations

Andrew reports that we got the sign of the tip correction wrong based on last night's data. So we haved changed the sign. Tonight we are using Tip=-15 Tilt=-5. This has the fortuitous effect of getting us off the -Y travel limit for M2.

The DX M2 Pointing Offsets are now dX=+0.255 dY=+0.766 dZ=0 RX=-147.3 RY=+49.1

Details

Open

01:42 Reset Shell

01:46 Kill and restart OSS to unstick M2 hexapod. (The hexapod was not power cycled.)

Assistant telescope operator Miller is starting up the telescope this evening. So John is enjoying the armchair operator's position.

02:04 Reset AZ encoders at 120.

02:10 Runa reports the wind on the balcony is 8 m/s (and nearly 0 in the chamber...ha ha)

02:12 Opening the shutter doors

02:23 pointto Pollux TelMode = TRACK gs=0 PAM-

02:28 tweak IE and CA to center star (high elevation). Put out of focus and remove coma by moving M2 RY=-200 RY=-40 and IE=-63 CA=-26

02:38 Runa report 5 m/s in chamber and 7 m/s on balcony

02:45 Take a cube of out of focus images for Runa irtc.20100405.024239.fits (single) and irtc.20100405_00079_cube.fits

02:46 pointto M5_0460 (low elevation) to set IA=-406 and IE=-75

02:52 Turned on Mirror Ventilation (forgot to when we opened)

02:54 pointto BS9143 ACQUIRE gs=0 PAM- to set IE=-63 CA=-23

02:59 pointto BS9143 ACTIVE gs=0 Collimate down to 200 nm rms in about 5 iterations

(must have missed some presets here in the log during the Doug-Fuyan transition)

Transformation Data

03:14 cl < home$tran_StoneE_00.cl seeing is 0.7-0.8 (prepdir before cl < home$tran_StoneE_00.cl )
star fits file cube file
PA 0
15 irtc.20100405.031533.fits irtc.20100405_00080_cube.fits
27 irtc.20100405.031748.fits irtc.20100405_00081_cube.fits
29 irtc.20100405.031954.fits irtc.20100405_00082_cube.fits
30 irtc.20100405.032202.fits irtc.20100405_00083_cube.fits
14 irtc.20100405.032413.fits irtc.20100405_00084_cube.fits
13 irtc.20100405.032623.fits irtc.20100405_00085_cube.fits
12 irtc.20100405.032837.fits irtc.20100405_00086_cube.fits
1 irtc.20100405.033048.fits irtc.20100405_00087_cube.fits

cl < tran_StoneE_mix1.cl
PA 90
27 irtc.20100405.033511.fits irtc.20100405_00088_cube.fits
29 irtc.20100405.033717.fits irtc.20100405_00089_cube.fits
30 irtc.20100405.033925.fits irtc.20100405_00090_cube.fits
PA 180
20 irtc.20100405.034146.fits irtc.20100405_00091_cube.fits
18 irtc.20100405.034359.fits irtc.20100405_00092_cube.fits
11 irtc.20100405.034606.fits irtc.20100405_00093_cube.fits
10 irtc.20100405.034813.fits irtc.20100405_00094_cube.fits
PA 270
11 irtc.20100405.035039.fits irtc.20100405_00095_cube.fits
10 irtc.20100405.035249.fits irtc.20100405_00096_cube.fits
9 irtc.20100405.035457.fits irtc.20100405_00097_cube.fits
7 irtc.20100405.035705.fits irtc.20100405_00098_cube.fits

03:58 cl < tran_StoneE_mix2.cl, during this script the seeing is at the range of 0.8-1.1.

PA 0
14 irtc.20100405.040029.fits irtc.20100405_00099_cube.fits
13 irtc.20100405.040235.fits irtc.20100405_00100_cube.fits
| 12 | irtc.20100405.040446.fits | irtc.20100405_00101_cube.fits | a faint nearby star 3" away
1 irtc.20100405.040654.fits irtc.20100405_00102_cube.fits
PA 90
28 irtc.20100405.040912.fits irtc.20100405_00103_cube.fits
23 irtc.20100405.041121.fits irtc.20100405_00104_cube.fits
24 irtc.20100405.041334.fits irtc.20100405_00105_cube.fits There is a bright star on the guider, guide on a wrong star
PA 180
9 irtc.20100405.041556.fits irtc.20100405_00106_cube.fits
7 irtc.20100405.041804.fits irtc.20100405_00107_cube.fits
23 irtc.20100405.042019.fits irtc.20100405_00108_cube.fits may be bad because the primary hits the limit,
24 irtc.20100405.042230.fits irtc.20100405_00109_cube.fits There is a bright star on the guider, guide on a wrong star
PA 270
8 irtc.20100405.042453.fits irtc.20100405_00110_cube.fits
6 irtc.20100405.042702.fits irtc.20100405_00111_cube.fits
17 irtc.20100405.042917.fits irtc.20100405_00112_cube.fits

04:30 cl < home$tran_StoneE_mix3.cl
PA 0
15 irtc.20100405.043221.fits irtc.20100405_00113_cube.fits
1 irtc.20100405.043434.fits irtc.20100405_00114_cube.fits
PA 90
21 irtc.20100405.043653.fits irtc.20100405_00115_cube.fits
22 irtc.20100405.043910.fits irtc.20100405_00116_cube.fits too dim
19 irtc.20100405.044130.fits irtc.20100405_00117_cube.fits
PA 180
21 irtc.20100405.044353.fits irtc.20100405_00118_cube.fits
22 irtc.20100405.044529.fits irtc.20100405_00119_cube.fits out of probe range. bad irtc image, background
19 irtc.20100405.044736.fits irtc.20100405_00120_cube.fits

primary against a limit, image may not be good
PA 270
5 irtc.20100405.045016.fits irtc.20100405_00121_cube.fits no star found in acquisition image
13 irtc.20100405.045229.fits irtc.20100405_00122_cube.fits primary against a limit, image may not be good
2 irtc.20100405.045447.fits rtc.20100405_00123_cube.fits There is a bright star on the guider, guide on a wrong star

05:05 pointto StoneIagw TelMode=ACQUIRE gs=1 PAM- change secondary RX to -300 and remove corrections of active optics in primary and click collimate on primary mirror. IE changes to -45

05:08 pointo StoneIagw TelMode=ACTIVE gs=0

05: 16 cl < home$tran_StoneI_mix1.cl
PA 45
18 irtc.20100405.051642.fits irtc.20100405_00124_cube.fits
21 irtc.20100405.051853.fits irtc.20100405_00125_cube.fits
25 irtc.20100405.052103.fits irtc.20100405_00126_cube.fits
PA 135
8 irtc.20100405.052326.fits irtc.20100405_00127_cube.fits
11 irtc.20100405.052541.fits irtc.20100405_00128_cube.fits
21 irtc.20100405.052757.fits irtc.20100405_00129_cube.fits
PA 225
10 irtc.20100405.053023.fits irtc.20100405_00130_cube.fits
14 irtc.20100405.053238.fits irtc.20100405_00131_cube.fits
5 irtc.20100405.053451.fits irtc.20100405_00132_cube.fits
PA 315
24 irtc.20100405.053755.fits irtc.20100405_00133_cube.fits
13 irtc.20100405.054009.fits irtc.20100405_00134_cube.fits
4 irtc.20100405.054223.fits irtc.20100405_00135_cube.fits

05:44 cl < home$tran_StoneI_mix2.cl
PA 225
11 irtc.20100405.054459.fits irtc.20100405_00136_cube.fits
8 irtc.20100405.054714.fits irtc.20100405_00137_cube.fits
13 irtc.20100405.054933.fits irtc.20100405_00138_cube.fits
6 irtc.20100405.055147.fits irtc.20100405_00139_cube.fits
3 irtc.20100405.055400.fits irtc.20100405_00140_cube.fits
PA 135
10 irtc.20100405.055626.fits irtc.20100405_00141_cube.fits
17 irtc.20100405.055840.fits irtc.20100405_00142_cube.fits
19 irtc.20100405.060059.fits irtc.20100405_00143_cube.fits
26 irtc.20100405.060312.fits irtc.20100405_00144_cube.fits
PA 45
19 irtc.20100405.060615.fits irtc.20100405_00145_cube.fits
22 irtc.20100405.060834.fits irtc.20100405_00146_cube.fits
PA 315
22 irtc.20100405.061059.fits irtc.20100405_00147_cube.fits
12 irtc.20100405.061310.fits irtc.20100405_00148_cube.fits
5 irtc.20100405.061527.fits irtc.20100405_00149_cube.fits
2 irtc.20100405.061740.fits irtc.20100405_00150_cube.fits

06:24 cl < home$tran_StoneI_mix3.cl
PA 315
25 irtc.20100405.062453.fits irtc.20100405_00151_cube.fits
14 irtc.20100405.062702.fits irtc.20100405_00152_cube.fits
6 irtc.20100405.062915.fits irtc.20100405_00153_cube.fits
3 irtc.20100405.063121.fits irtc.20100405_00154_cube.fits
PA 225
9 irtc.20100405.063343.fits irtc.20100405_00155_cube.fits
16 irtc.20100405.063555.fits irtc.20100405_00156_cube.fits
12 irtc.20100405.063808.fits irtc.20100405_00157_cube.fits
4 irtc.20100405.064020.fits irtc.20100405_00158_cube.fits
2 irtc.20100405.064230.fits irtc.20100405_00159_cube.fits
PA 135
9 irtc.20100405.064453.fits irtc.20100405_00160_cube.fits
16 irtc.20100405.064706.fits irtc.20100405_00161_cube.fits
18 irtc.20100405.064919.fits irtc.20100405_00162_cube.fits
20 irtc.20100405.065132.fits irtc.20100405_00163_cube.fits
27 irtc.20100405.065340.fits irtc.20100405_00164_cube.fits
PA 45
20 irtc.20100405.065640.fits irtc.20100405_00165_cube.fits
26 irtc.20100405.065852.fits irtc.20100405_00166_cube.fits

06:59 offset 0 -40 to take the back ground image. Then the shell is dead, reset.

07:39 pointto BS9152 TelMode=ACQUIRE gs=0 PAM- The elevation = 69
pointto BS9177 TelMode=ACQUIRE gs=0 PAM- The elevation = 31
Primary against the limit.
start collimate file: wfsc278
Shell reset.

08:30 pointto BS9152 TelMode=ACQUIRE gs=0 PAM-.
pointto BS9152 TelMode=ACTIVE gs=0,
active mode error.
restart GCSR.
resend the preset.. Primary hits the limit, change RX in secondary from -300 to -400. remove the correction of active optics in primary. resend the preset. lost the guide star. go to the acquire mode and set RX to -350 and adjust IE and CA to put the star in the center and then set RX to -400 and play with IE and CA to put the star in the center. Go to active mode to collimate.

Field Aberration Data

08:54 cl < home$field_BS9152.cl
The primary hits the limit. Set RX to 420. The wrong script name uses LFBG parameters.

09:20 cl < home$newfield_BS9152_right.cl

GStar wfscimage#
Sequence 1Sorted ascending
Sequence 2
0 9 - 13
17 14 - 16
0 17 - 19
18 20 - 22
0 23 - 25
7 26 - 28
0 29 - 31
4 32 - 34
0 35 - 37
1 38 - 40
0 41 - 43
9 44 - 46
0 47 - 49
0 50 - 54
17 55 - 57
0 58 - 60
18 61 - 63
0 64 - 66
7 67 - 69
0 70 - 72
4 73 - 75
0 76 - 78
1 79 - 81
0 82 - 84
9 85 - 87
0 88 - 90

10:00 Discover that right side instrument cooling loop temperature is up to 7 degC when it should be 0 degC. From Runa's telemetry plot, it seems the failure happened about midnight (3 hours ago). Reading on IC page of ECS says CV0424 is at 103%. But when John sends a manual command to 50%, the apparent value jumps to 2% (closed). Note that there is no ECS alarm for this case (pumps are still running). See Issue 2648.

Low-elevation collimation

11:02 pointto BS9131 TelMode=ACTIVE gs=0 elevation=22 to check collimation at low elevation.
Apr  5 11:04:24 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 588.573 0 0 -110.879 537.464 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:05:08 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 473.947 631.846 606.183 -52.7452 366.097 108.491 -17.2522 -248.401 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:06:03 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -851.171 0 0 68.2029 125.179 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:06:43 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -578.142 920.859 99.5814 7.29454 69.8581 -228.593 -5.74796 164.425 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:07:24 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -275.269 436.574 -28.0351 29.808 -3.72734 -134.092 -4.71848 65.4037 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:08:03 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -298.161 444.686 90.6393 -24.5791 97.7588 -176.521 74.0611 89.8753 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:08:44 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -366.152 362.105 16.0823 11.1312 -5.89536 -137.61 17.8443 75.3973 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:09:23 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -209.034 287.249 136.524 4.68573 46.312 -6.53692 26.0713 25.0467 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:10:03 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -123.493 185.021 59.6372 20.8754 10.384 -51.0525 13.4202 31.1355 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:10:43 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -36.0132 148.44 30.1374 -10.1305 -22.0116 -76.276 2.42176 2.37393 -0 -0 -0 -0 -0 -0 0 -0 0 -0 -14.9297
Apr  5 11:11:23 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -180.934 259.293 79.4949 9.39771 17.759 -54.3403 16.057 19.9945 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:12:02 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -150.858 217.308 82.6865 6.67482 3.04474 -56.5111 31.5071 52.4288 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:12:41 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -81.5015 146.531 91.5999 21.6089 5.18267 -65.6098 27.9636 26.633 0 0 0 0 0 0 0 0 0 0 0
Apr  5 11:13:21 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 124.127 60.3453 10.544 -33.0009 20.6949 -10.5354 32.5323 13.2056 -0 -0 -0 -0 -0 0 0 -0 0 -0 -19.1628
Apr  5 11:14:02 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -12.1108 129.349 100.722 -2.78523 31.9094 -36.9729 39.3651 29.8172 -0 -0 -0 -0 -0 -0 0 -0 0 -0 -18.0617
Apr  5 11:14:41 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -78.6121 112.729 23.2599 -65.3618 8.25143 -17.293 35.1441 1.42131 -0 -0 -0 -0 -0 0 0 -0 0 -0 -12.5818
Apr  5 11:15:22 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -68.7581 30.7959 77.2477 22.0484 25.0274 -30.5641 52.9028 5.20929 -0 0 -0 -0 -0 -0 -0 -0 0 -0 -14.0696
Apr  5 11:16:02 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -16.924 82.1336 -70.8436 10.3042 6.38059 -14.7091 18.2819 -16.6667 -0 -0 -0 -0 -0 -0 0 -0 0 -0 -10.6517
Apr  5 11:16:41 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 57.7932 -26.0198 31.104 22.3348 -12.929 -47.7966 -10.3355 -13.8259 -0 0 -0 -0 -0 -0 0 -0 0 -0 -12.6909
Apr  5 11:17:21 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 -17.0492 -62.611 37.4457 32.4732 -17.8366 -36.4707 -1.46401 -21.4575 -0 0 -0 -0 -0 -0 0 -0 0 -0 -8.35444
Apr  5 11:18:01 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 89.2932 98.9705 9.51283 -7.46258 7.4654 -15.1122 23.8037 -20.636 -0 -0 -0 -0 -0 0 0 -0 0 -0 -12.036
Apr  5 11:18:41 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 177.75 -11.7824 -5.29433 28.6938 3.08855 -13.4451 -16.8335 -6.17109 -0 0 -0 -0 -0 -0 0 -0 0 -0 -16.4399

11:20 The DX primary mirror is panic, raise the primary and there is a big coma. Why was this coma vastly different after we recovered the panic?

M3 Tests - Repeatability

11:38 pointto BS9177 TelMode=ACQUIRE gs=0 PAM-
center the guide star in the acquire field,

11:40 Test: home the Tertiary mirror the guide star disappear, and then collimate Tertiary mirror the star coming back within ~1 arcsec (limited by tracking drift)

11:48 Test: put tip and tilt -200 and 200, home the Tertiary mirror, take off the values and collimate Tertiary mirror the star coming back within ~1 arcsec.

11:53 Test: Home the Tertiary mirror selector and go back to RFBG, the star coming back again within ~1 arcsec.

M3 Tests - Scale of Motion

move the star from the center to the edge of the guider acquire field. First the tip = +140 arcsec (moves to low on left edge), and then the tilt = +230 arcsec (moves to top near upper left corner). This seems like correct behaviour as far as our sleepy brains can tell. Coordinates are rotated from parallactic by the selector angle to feed the particular focal station.

12:05 pointto BS9177 TelMode=ACTIVE, gs=0 .
GCS is dead (hung on the preset), restart GCSR and resend the preset.

Induced Panic

12:12 Press the button to induce an artificial panic to see of the collimation position is repeatable.

12:20 Torque switch set in hardpoint 0. This ends our night with DX M1.

Close

12:30 Close Dome

Topic revision: r6 - 29 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback